Sahil makes car payments of $318/month for the next 5 years.


His car loan has an interest rate of 2.6%, discounted monthly.


What was the initial price of the car?



1
Expert's answer
2021-04-29T17:11:56-0400

"PV=P\times\frac{1-(1+r)^{-n}}{r}"

PV = present value

P = value of each payment

r = interest rate per period

n = number of periods


"PV=318\times \frac{1-(1+0.026)^{-60}}{0.026}"


"=318 \times 30.21664581"

=$ "9,608.893368"